# Minimizing Human Intervention in Online Classification

### v1

@@ 4.2 Center-based Classifier (CC)
Let $p_{\min}:=\min_{i\in[N]}p_{i}.$ Under Assumptions 3.2(ii) and 3.3(ii), if $T\leq e^{d},$ the regret of CC satisfies
@@ 5 NUMERICAL EXPERIMENTS
In this section, we evaluate GHC on a real-world dataset using various embedding models. The reward values are instantiated as $\alpha=-1$, $\beta=+1$, $\gamma=-10$. Additional experiments on both synthetic and real-world datasets are deferred to Appendix F, where we notably show that GHC outperforms both CC and a sequential $k$-means baseline.
To evaluate GHC in a realistic setting, we introduce the Quora Question Group (QQG) dataset. Each question in the dataset is assigned one of $N=1103$ labels, where each label corresponds to a group of questions that can be addressed by a single answer. We compare the performance of three different retrievers/embedding models: Nomic333huggingface.co/nomic-ai/nomic-embed-text-v1 (Nussbaum et al.,, 2024), E5444huggingface.co/intfloat/e5-large(Wang et al.,, 2022) and Mistral_E5 555huggingface.co/intfloat/e5-mistral-7b-instruct (Wang et al.,, 2024) (see Appendix F.2.2). We first encode each question with the chosen retrievers, producing embeddings of dimension 784 for Nomic, 1,024 for E5, and 4,096 for Mistral_E5. Given the relatively small size of our datasets (each label corresponds to a few questions only), we initialize the algorithm by picking an example question for each label, thereby obviating the first phase of GHC (where CHC is applied until every label is represented).
Fig. 4(a) shows the average accumulated regret for the three models when we use their respective optimal threshold. Results are averaged over three independent runs. As expected, Mistral_E5, the most recent and largest model in our experiments with 7 billion parameters, outperforms both E5 (330 million parameters) and Nomic (110 million parameters). Fig. 4(b) shows each model’s performance across all threshold values $\tau$. Observe that the optimal thresholds are quite high, which is expected since our embeddings are high-dimensional.
In Appendix F.2.3, we extend our evaluation to the ComQA (Abujabal et al.,, 2019) and CQADupStack (Hoogeveen et al.,, 2015) datasets. ComQA contains open-domain questions derived from Wikipedia, whereas CQADupStack comprises expert-domain questions from technical forums, including topics such as physics and mathematics. We provide additional details about our datasets in Appendix F.2.1.
(a) Accumulated regret
@@
Figure 4: Comparison of GHC using different LLMs on Quora Question Groups dataset.
@@ 6 CONCLUSION
We introduced the problem of online classification with expert guidance and proposed algorithms tailored to different budget regimes. In particular, we developed the CC and CHC algorithms, which enjoy strong regret guarantees in low- and high-horizon settings, respectively, and further proposed the more versatile GHC algorithm, which demonstrates strong empirical performance across all regimes. A natural next step is to provide a theoretical regret analysis of this latter algorithm.
More broadly, the online learning problem studied here represents a first instance of a broader class of challenges aimed at minimizing costly human intervention during model training or fine-tuning. We anticipate that many more such problems will emerge, especially in the context of fine-tuning or adapting large foundation models using human feedback.
@@
      An analysis of the properties and complexity (time, space, sample size) of any algorithm. [Yes] The algorithm’s computational complexities are discussed in Sections 4 and 4.3. A regret analysis of CHC is also provided.
@@
      Citations of the creator If your work uses existing assets. [Yes] The prior works that introduced the datasets or LLMs we use for our numerical experiments are all cited in the paper and direct links are provided.
@@ Appendix B PROOFS OF SECTION 4
We rely on the following Theorem from Bárány and Buchta, (1993).
Proposition B.2 (Theorem 2 from Bárány and Buchta, (1993)).
@@
To derive the regret bound in the spherical setting, we first obtain a variant of Corollary B.3 by projecting spherical convex polytopes to Euclidean polytopes via the gnomonic projection (see e.g. Besau and Schuster, (2016)), similarly to Besau et al., (2018) in the proof of their Theorem 1.3. This result holds for distributions with bounded densities w.r.t. the spherical Lebesgue measure $\omega.$
@@
denote the gnomonic projection from the open halfsphere $\mathcal{S}_{e}^{+}$ onto the hyperplane tangent to $\mathcal{S}^{d-1}$ at the pole $e$ of $\mathcal{S}_{e}^{+}$. This projection is bijective and its inverse is given by $g_{e}^{-1}(y)=(y+e)/\|y+e\|_{2}$ (Besau et al.,, 2018). Let $g_{e}(P). QQ$ is a Euclidean convex polytope in $\mathbb{R}^{d-1}$. Let $\rho=g_{e}\#\nu$ denote the pushforward measure of $\nu$ onto $Q$, i.e., $\rho(A)=\nu(g_{e}^{-1}(A))$ for all measurable set $A\subseteq Q.$ $\rho$ is a probability measure on $Q$. Let $q_{i}=g_{e}(p_{i})$ for $i=1,\dots,n$. The points $q_{i}$ are independent samples from $\rho$. Let $Q_{n}=\mathrm{conv}(q_{1},\dots,q_{n})$ be their Euclidean convex hull.
@@
We now verify the conditions of Corollary B.3 for our measure $\rho$ on $Q$. By assumption, $\nu$ has a density $f_{P}=\mathrm{d}\nu/\mathrm{d}\omega$ w.r.t. the spherical Lebesgue measure $\omega$ satisfying $0<c\leq f_{P}(y)\leq C$ for $\omega$-a.e. every $y\in P.$ The pushforward measure $\rho=g_{e}\#\nu$ has a density $f_{Q}=d\rho/d\lambda$ on $Q$ given by $f_{Q}(x)\propto f_{P}(g_{e}^{-1}(x))(1+\|x\|^{2})^{-d/2}$ for $x\in Q$ (see Besau et al., (2018), p.36). On the compact $Q$, the density $f_{Q}$ is bounded below by $c^{\prime}=\min_{x\in Q}f_{Q}(x)>0$ and above by $C^{\prime}=\max_{x\in Q}f_{Q}(x)<\infty$, where $\frac{C^{\prime}}{c^{\prime}}=\frac{C}{c}\left(\frac{1+r^{2}_{\max}}{1+r^{2}_{\min}}\right)^{d/2}$ for $r_{\max}=\max_{x\in g_{e}(P)}\|x\|$, $r_{\min}=\min_{x\in g_{e}(P)}\|x\|$. Note that for $x=g_{e}(y)\in g_{e}(P)$ , $\|x\|^{2}=\|\frac{y}{y^{\top}e}-e\|^{2}=\frac{1}{(y^{\top}e)^{2}}-1$ since $y,e\in\mathcal{S}^{d-1}.$ Thus
@@
We first show the following result, which is essentially a dual version of the upper bound from Section 2 of Besau et al., (2018).
@@
Consider the dual polytope $P^{o}$ of $P$, which has $k$ vertices and the same number of flags as $P$. By the upper bound theorem (McMullen, (1970), or Theorem 6.6 in Billera and Ehrenborg, (2000)), $F(P^{o})\leq F(C_{d}(k))$ where $C_{d}(k)$ is the cyclic polytope in $\mathbb{R}^{d}$ with $k$ vertices. Since $C_{d}(k)$ is simplicial, $F(C_{d}(k))=d!f_{d-1}(C_{d}(k))$, where $f_{d-1}(C_{d}(k))$ is the number of (d1)(d-1)-dimensional faces of $C_{d}(k)$. Hence, we have
@@
(see Ziegler, (1998), p.25). This concludes the proof. ∎
@@
When $\mathcal{E}=\mathcal{S}^{d-1}$, the argument is slightly simpler since only the N1N-1 boundaries of the Voronoi tessellation can form facets of the cells (the space $\mathcal{S}^{d-1}$ does not have boundaries to consider that might increase their number). More precisely, $\mathcal{C}_{i}$ is a spherically convex polytope formed by at most N1N-1 great spheres of $\mathcal{E}$, so it has at most N1N-1 faces of dimension d2d-2. We first remark that under the open halfsphere assumption, each cell has at least $d$ facets (Theorem 6.3.7 in Ratcliffe, (2019)). Consequently, we must have $N\geq d+1$. A spherical analogue of Lemma B.7 applies and it gives
@@
In this section, we present a more general version of Theorem B.13 along with its proof. We recall that the expert labeling policy is given by the Voronoi tessellation with seeds $s_{i}$, as described in Section 3.2. We show that in the regime $T\leq e^{d}$, if the queries are distributed as a subgaussian mixture with sufficiently separated centers, a simple Center-based Classifier (CC) can achieve $N\log{N}$ regret after $T$ rounds when the mixture is homogeneous. CC’s general regret bound holds under the following condition, which implies 3.3(ii).
@@ B.2.2 Bounding the Second Phase Regret
In the proof of Proposition B.16, we used the simple upper bound $\mathbb{E}[\tau_{1}]\leq(\log{N}+1)/p_{\min}.$ $\mathbb{E}[\tau_{1}]$ can in fact be computed exactly as $I:=\int_{0}^{\infty}\left(1-\prod_{i=1}^{N}(1-e^{-p_{i}t})\right)\mathrm{d}t$ (see example 5.17 in Ross, (2010)). Using this formula would yield a sharper but less readable regret bound that scales with $I$ instead of $(\log{N})/p_{\min}$ in the $T\leq e^{d}$ regime. We also note that an alternative bound is $I\leq\sum_{i=1}^{N}1/p_{i},$ which can be sharper than $I\leq(\log{N}+1)/p_{\min}$ depending on the value of the weights $p_{i}.$
@@ Appendix D ADDITIONAL RELATED WORK
Our theoretical analysis relies on results regarding the approximation of convex bodies by random polytopes. We refer to the recent survey of Prochno et al., (2022), and to the books of Ziegler, (1998) and Ratcliffe, (2019) for general background on Euclidean and spherical geometry.
@@
Our problem formulation (Section 3) assumes queries are represented by embeddings, and our real-world experiments (Section 5, Appendix F.2.3) utilize state-of-the-art embedding models derived from LLMs. Such embeddings are foundational to many modern Natural Language Processing (NLP) applications: for example, in contemporary applications of NLP to real-world question-answering environments, where knowledge may reside in documentation, Retrieval-Augmented Generation (RAG) (Lewis et al.,, 2020) has emerged as a pivotal technique. A RAG system comprises two main components: a Retriever and a Large Language Model. Retrievers excel at representing similar words and sentences closely in the embedding space, thereby understanding language patterns effectively. Notable retriever models, such as Dense Passage Retrieval (DPR) (Karpukhin et al.,, 2020), EmbEddings from bidirEctional Encoder rEpresentations (E5) (Wang et al.,, 2022), and General Text Embedding (GTE) (Li et al.,, 2023), leverage pretrained architectures such as BERT (Devlin et al.,, 2019) to initialize encoders. These encoders are fine-tuned to ensure that the cosine similarity between the query and passage accurately captures their true relationship.
Recently, the NLP community has increasingly favored decoder architectures for creating embeddings (Springer et al.,, 2024; BehnamGhader et al.,, 2024), as these approaches have demonstrated superior performance over traditional encoder-based methods. Among these, Mistral-E5 (Wang et al.,, 2024) stands out as the state-of-the-art LLM for text retrieval. Retriever models are commonly evaluated using benchmarks such as BEIR (Thakur et al.,, 2021) and MTEB (Muennighoff et al.,, 2022).
@@ E.2 Asymptotic Lower bound on the Regret of CHC
The lower bounds established in Theorem E.2, are independent of the distribution $\mu$ and match the upper bounds of Theorem 4.1 in the uniform case ($C/c=1$), up to a multiplicative factor linear in the dimension $d$. This suggests that CHC does not exploit favorable scenarios where $\mu$ is highly concentrated around the seed queries of the Voronoi tessellation. It is important to stress that this lower bound is specific to the CHC algorithm and does not extend to all possible algorithms. The proof, presented below, relies on the connection between random polytopes and the so-called floating body of the polytope they approximate, a notion developed in Schütt and Werner, (1990).
@@
Lemma E.3 leverages the relationship between the uncovered region $P\setminus P_{n}$ and the wet part of $\nu$, which we now define. For a distribution $\nu$, the wet part of $\nu$ is defined in Bárány et al., (2020) as
@@
To understand this notion intuitively, consider the case where $\nu$ is uniform on the unit ball, and assume that the ball is filled with a volume $t$ of water. The wet part $W^{t}_{\nu}$ represents the “trace” left by the water as the ball rolls on a flat surface, and $w^{\nu}(t)$ is the volume of that wet part. The $\nu$-measure of the wet part is tightly linked to the expected missing volume $\mathbb{E}[1-\nu(P_{n})]$, as shown in e.g. Schütt, (1991) and Bárány and Buchta, (1993) for the case where $\nu$ is uniform on some convex polytope of $\mathbb{R}^{d}$, and in Bárány et al., (2020) for arbitrary measures in $\mathbb{R}^{d}.$ Such results are sometimes stated in terms of the floating body of $\nu$, which is the relative complement of its wet part in its support, see Schütt and Werner, (1990).
@@
By the lower bound of Theorem 1.2 in Bárány et al., (2020):
@@
This lower bound is straightfoward to derive: one can note that $\mathbb{P}(x\notin P_{n})\geq(1-t)^{n}$ for any $x\in W^{t}_{\nu}$, thus $\mathbb{E}[\nu(P\setminus P_{n})]=\int_{0}^{\infty}\mathbb{P}(x\notin P_{n})d\mu(x)\geq(1-t)^{n}w^{\nu}(t)$. Setting $t=1/n$ yields the result (see Section 3.1 in Bárány et al., (2020) for a more detailed proof).
Then, by Corollary 1.2 in Besau et al., (2018), under Assumption 3.3(ii), the asymptotic behavior of $w^{\nu}$ is independent of $\nu$:
@@
    The geometry of each cell $\mathcal{C}_{i}$. This dependency comes from our application of Proposition B.2 (Theorem 2 from Bárány and Buchta, (1993)).
@@
Below, we support the claim that it is very difficult to identify the precise dependency in the cell geometry (and in particular in the dimension) induced by Theorem $2$ from Bárány and Buchta, (1993), which was not tracked in their original proof. One way to attempt to identify this dependency is to exploit the relationship between random polytopes and the floating body of the polytope they approximate, as already leveraged in the proof of Theorem E.2. For example, Bárány et al., (2020) show in their Theorem 1.2 that if $p_{1},\dots,p_{n}$ are sampled according to some arbitrary $\nu$ on $\mathbb{R}^{d}$ with convex hull $P_{n}$,
@@
Here, the dependency in $(\log{n})/n$ in the wet part measure is necessary because $\nu$ is arbitrary, but for more specific distributions (e.g. log-concave), this dependency may potentially be improved to $1/n$, as discussed in Section 2.4 of Bárány et al., (2020). Since $\mathbb{E}[1-\nu(P_{n})]$ is precisely the measure of the region uncovered by $P_{n}$, we may attempt to upper bound $w^{\nu}(t)$ for fixed $t$ to obtain a fully explicit upper bound on our quantity of interest. This is not too difficult when $\nu$ is uniformly distributed on the hypercube $\mathcal{I}^{d}$, as we show below.
@@
Let $v(x)=\min\{\lambda(\mathcal{I}^{d}\cap H),x\in H\text{ and $H$ is a halfspace}\}$, and $u(x)=\lambda(\mathcal{I}^{d}\setminus(2x-\mathcal{I}^{d}))$. This is the volume of a Macbeath region, see e.g. Bárány, (2008), where they recall in their Lemma 4.3 that $u(x)\leq 2v(x)$. Note also that the wet part for the uniform distribution on $\mathcal{I}^{d}$ is $\{x\in K,v(x)\leq t\lambda(\mathcal{I}^{d})\}.$ Then

### v2

@@ 3.2 Center-based Classifier (CC)
Let $p_{\min}:=\min_{i\in[N]}p_{i}.$ Under Assumptions 2.2(ii) and 2.3(ii), if $T\leq e^{d},$ the regret of CC satisfies
@@ 4 NUMERICAL EXPERIMENTS
In this section, we evaluate GHC on a real-world dataset using various embedding models, and compare GHC with CC and baselines from the literature. The reward values are instantiated as $\alpha=-1$, $\beta=+1$, $\gamma=-10$. Additional experiments on both synthetic and real-world datasets are deferred to Appendix F. The code used for our experiments is available at github.com/wilrev/MinimizingHumanIntervention.
To evaluate GHC in a realistic setting, we introduce the Quora Question Group (QQG) dataset, which is constructed from the Quora Question Pairs dataset (Wang et al., 2017). Each question in the dataset is assigned one of $N=1103$ labels, where each label corresponds to a group of questions that can be addressed by a single answer. We compare the performance of three different retrievers/embedding models: Nomic333huggingface.co/nomic-ai/nomic-embed-text-v1 (Nussbaum et al., 2024), E5444huggingface.co/intfloat/e5-large(Wang et al., 2022) and Mistral_E5 555huggingface.co/intfloat/e5-mistral-7b-instruct (Wang et al., 2024) (see Appendix F.2.2). We first encode each question with the chosen retrievers, producing embeddings of dimension 784 for Nomic, 1,024 for E5, and 4,096 for Mistral_E5. Given the relatively small size of our datasets (each label corresponds to a few questions only), we initialize the algorithm by picking an example question for each label, thereby obviating the first phase of GHC (where CHC is applied until every label is represented).
Fig. 4(a) shows the average cumulative regret for the three models when we use their respective best performing threshold. Results are averaged over three independent runs. As expected, Mistral_E5, the most recent and largest model in our experiments with 7 billion parameters, outperforms both E5 (330 million parameters) and Nomic (110 million parameters). Fig. 4(b) shows each model’s performance across all threshold values $\tau$. We observe that the best performing thresholds are quite high, which is expected since our embeddings are high-dimensional.
In Appendix F.2.3, we extend our evaluation to the ComQA (Abujabal et al., 2019) and CQADupStack (Hoogeveen et al., 2015) datasets. ComQA contains open-domain questions derived from Wikipedia, whereas CQADupStack comprises expert-domain questions from technical forums, including topics such as physics and mathematics. We provide additional details about the datasets used in our experiments in Appendix F.2.1.
(a) Regret vs. step ($\tau=0.9$)
@@
Figure 4: Comparison of GHC using different text embedding models on Quora Question Groups dataset.
We then compare the performance of GHC, CC (Section 3.2) and two baselines on the QQG dataset, using embeddings generated by Mistral_E5. The first baseline (SKM) is based on sequential $k$-means (MacQueen, 1967). After a supervised initialization phase, this algorithm updates centroids at each round and always guesses based on the closest centroid. The second baseline (AMP) is an active variant of the multiclass perceptron (Crammer and Singer, 2003; Cesa-bianchi et al., 2004). It defers to the expert when its confidence margin falls below a tunable threshold $\tau\in[0,1].$ Both baselines are described in more detail in Appendix F.1.1. In Fig. 5, we report the average cumulative regret of each algorithm over three independent runs, selecting the best performing threshold for GHC and AMP among a range of thresholds (see Appendix F.2.4 for the specific ranges tested). CC exhibits approximately linear regret, which is expected since $\log{N}/p_{\min}$ is not negligible with respect to the horizon $T$ in this experiment (see first paragraph of Section 3.3). The unsupervised baseline SKM also exhibits similar regret scaling, which indicates that both algorithms fail to effectively learn the data structure over time. In contrast, both GHC and AMP exhibit sublinear regret. GHC consistently achieves the lowest cumulative regret, and its performance gap with AMP widens over time, highlighting the superior long-term learning capability of GHC. Additional comparisons other datasets and with different embedding models are deferred to Appendix F.2.4.
(a) Regret vs. step
Figure 5: Comparison of different Algorithms on Quora Question Groups dataset with Mistral_E5.
@@ 5 CONCLUSION
We proposed algorithms tailored to different budget regimes to address the problem of online classification with expert guidance. In particular, we developed the CC and CHC algorithms, which enjoy strong regret guarantees in low- and high-horizon settings, respectively, and further proposed the more versatile GHC algorithm, which demonstrates strong empirical performance across all regimes. A natural next step is to provide a theoretical regret analysis of this latter algorithm.
More broadly, the online learning problem studied here represents an instance of a broader class of challenges aimed at minimizing costly human intervention during model training or fine-tuning. We anticipate that many more such problems will emerge, especially in the context of fine-tuning or adapting large foundation models using human feedback.
6 ACKNOWLEDGMENTS
This work was supported by computational resources from the National Academic Infrastructure for Supercomputing in Sweden (NAISS), funded in part by the Swedish Research Council (grant no. 2025/5-542).
@@ 6 ACKNOWLEDGMENTS
      An analysis of the properties and complexity (time, space, sample size) of any algorithm. [Yes] The algorithm’s computational complexities are discussed in Sections 3 and $3.$ A regret analysis of CHC is also provided.
@@
      Citations of the creator If your work uses existing assets. [Yes] The prior works that introduced the datasets or embedding models we use for our numerical experiments are all cited in the paper and direct links are provided.
@@ Appendix B PROOFS OF SECTION 3
We rely on the following Theorem from Bárány and Buchta (1993).
Proposition B.2 (Theorem 2 from Bárány and Buchta (1993)).
@@
To derive the regret bound in the spherical setting, we first obtain a variant of Corollary B.3 by projecting spherical convex polytopes to Euclidean polytopes via the gnomonic projection (see e.g. Besau and Schuster (2016)), similarly to Besau et al. (2018) in the proof of their Theorem 1.3. This result holds for distributions with bounded densities w.r.t. the spherical Lebesgue measure $\omega.$
@@
denote the gnomonic projection from the open halfsphere $\mathcal{S}_{e}^{+}$ onto the hyperplane tangent to $\mathcal{S}^{d-1}$ at the pole $e$ of $\mathcal{S}_{e}^{+}$. This projection is bijective and its inverse is given by $g_{e}^{-1}(y)=(y+e)/\|y+e\|_{2}$ (Besau et al., 2018). Let $g_{e}(P). QQ$ is a Euclidean convex polytope in $\mathbb{R}^{d-1}$. Let $\rho=g_{e}\#\nu$ denote the pushforward measure of $\nu$ onto $Q$, i.e., $\rho(A)=\nu(g_{e}^{-1}(A))$ for all measurable set $A\subseteq Q.$ $\rho$ is a probability measure on $Q$. Let $q_{i}=g_{e}(p_{i})$ for $i=1,\dots,n$. The points $q_{i}$ are independent samples from $\rho$. Let $Q_{n}=\mathrm{conv}(q_{1},\dots,q_{n})$ be their Euclidean convex hull.
@@
We now verify the conditions of Corollary B.3 for our measure $\rho$ on $Q$. By assumption, $\nu$ has a density $f_{P}=\mathrm{d}\nu/\mathrm{d}\omega$ w.r.t. the spherical Lebesgue measure $\omega$ satisfying $0<c\leq f_{P}(y)\leq C$ for $\omega$-a.e. every $y\in P.$ The pushforward measure $\rho=g_{e}\#\nu$ has a density $f_{Q}=d\rho/d\lambda$ on $Q$ given by $f_{Q}(x)\propto f_{P}(g_{e}^{-1}(x))(1+\|x\|^{2})^{-d/2}$ for $x\in Q$ (see Besau et al. (2018), p.36). On the compact $Q$, the density $f_{Q}$ is bounded below by $c^{\prime}=\min_{x\in Q}f_{Q}(x)>0$ and above by $C^{\prime}=\max_{x\in Q}f_{Q}(x)<\infty$, where $\frac{C^{\prime}}{c^{\prime}}=\frac{C}{c}\left(\frac{1+r^{2}_{\max}}{1+r^{2}_{\min}}\right)^{d/2}$ for $r_{\max}=\max_{x\in g_{e}(P)}\|x\|$, $r_{\min}=\min_{x\in g_{e}(P)}\|x\|$. Note that for $x=g_{e}(y)\in g_{e}(P)$ , $\|x\|^{2}=\|\frac{y}{y^{\top}e}-e\|^{2}=\frac{1}{(y^{\top}e)^{2}}-1$ since $y,e\in\mathcal{S}^{d-1}.$ Thus
@@
We first show the following result, which is essentially a dual version of the upper bound from Section 2 of Besau et al. (2018).
@@
Consider the dual polytope $P^{o}$ of $P$, which has $k$ vertices and the same number of flags as $P$. By the upper bound theorem (McMullen (1970), or Theorem 6.6 in Billera and Ehrenborg (2000)), $F(P^{o})\leq F(C_{d}(k))$ where $C_{d}(k)$ is the cyclic polytope in $\mathbb{R}^{d}$ with $k$ vertices. Since $C_{d}(k)$ is simplicial, $F(C_{d}(k))=d!f_{d-1}(C_{d}(k))$, where $f_{d-1}(C_{d}(k))$ is the number of (d1)(d-1)-dimensional faces of $C_{d}(k)$. Hence, we have
@@
(see Ziegler (1998), p.25). This concludes the proof. ∎
@@
When $\mathcal{E}=\mathcal{S}^{d-1}$, the argument is slightly simpler since only the N1N-1 boundaries of the Voronoi tessellation can form facets of the cells (the space $\mathcal{S}^{d-1}$ does not have boundaries to consider that might increase their number). More precisely, $\mathcal{C}_{i}$ is a spherically convex polytope formed by at most N1N-1 great spheres of $\mathcal{E}$, so it has at most N1N-1 faces of dimension d2d-2. We first remark that under the open halfsphere assumption, each cell has at least $d$ facets (Theorem 6.3.7 in Ratcliffe (2019)). Consequently, we must have $N\geq d+1$. A spherical analogue of Lemma B.7 applies and it gives
@@
In this section, we present a more general version of Theorem B.13 along with its proof. We recall that the expert labeling policy is given by the Voronoi tessellation with seeds $s_{i}$, as described in Section 2.2. We show that in the regime $T\leq e^{d}$, if the queries are distributed as a subgaussian mixture with sufficiently separated centers, a simple Center-based Classifier (CC) can achieve $N\log{N}$ regret after $T$ rounds when the mixture is homogeneous. CC’s general regret bound holds under the following condition, which implies 2.3(ii).
@@ B.2.2 Bounding the Second Phase Regret
In the proof of Proposition B.16, we used the simple upper bound $\mathbb{E}[\tau_{1}]\leq(\log{N}+1)/p_{\min}.$ $\mathbb{E}[\tau_{1}]$ can in fact be computed exactly as $I:=\int_{0}^{\infty}\left(1-\prod_{i=1}^{N}(1-e^{-p_{i}t})\right)\mathrm{d}t$ (see example 5.17 in Ross (2010)). Using this formula would yield a sharper but less readable regret bound that scales with $I$ instead of $(\log{N})/p_{\min}$ in the $T\leq e^{d}$ regime. We also note that an alternative bound is $I\leq\sum_{i=1}^{N}1/p_{i},$ which can be sharper than $I\leq(\log{N}+1)/p_{\min}$ depending on the value of the weights $p_{i}.$
@@ Appendix D ADDITIONAL RELATED WORK
Our theoretical analysis relies on results regarding the approximation of convex bodies by random polytopes. We refer to the recent survey of Prochno et al. (2022), and to the books of Ziegler (1998) and Ratcliffe (2019) for general background on Euclidean and spherical geometry.
@@
Our problem formulation (Section 2) assumes queries are represented by embeddings, and our real-world experiments (Section 4, Appendix F.2.3) utilize state-of-the-art text embedding models. Such embeddings are foundational to many modern Natural Language Processing (NLP) applications: for example, in contemporary applications of NLP to real-world question-answering environments, where knowledge may reside in documentation, Retrieval-Augmented Generation (RAG) (Lewis et al., 2020) has emerged as a pivotal technique. A RAG system comprises two main components: a Retriever and a Large Language Model. Retrievers excel at representing similar words and sentences closely in the embedding space, thereby understanding language patterns effectively. Notable retriever models, such as Dense Passage Retrieval (DPR) (Karpukhin et al., 2020), EmbEddings from bidirEctional Encoder rEpresentations (E5) (Wang et al., 2022), and General Text Embedding (GTE) (Li et al., 2023), leverage pretrained architectures such as BERT (Devlin et al., 2019) to initialize encoders. These encoders are fine-tuned to ensure that the cosine similarity between the query and passage accurately captures their true relationship.
Recently, the NLP community has increasingly favored decoder architectures for creating embeddings (Springer et al., 2024; BehnamGhader et al., 2024), as these approaches have demonstrated superior performance over traditional encoder-based methods. Among these, Mistral-E5 (Wang et al., 2024) stands out as the state-of-the-art embedding model for text retrieval. Retriever models are commonly evaluated using benchmarks such as BEIR (Thakur et al., 2021) and MTEB (Muennighoff et al., 2022).
@@ E.2 Asymptotic Lower bound on the Regret of CHC
The lower bounds established in Theorem E.2, are independent of the distribution $\mu$ and match the upper bounds of Theorem 3.1 in the uniform case ($C/c=1$), up to a multiplicative factor linear in the dimension $d$. This suggests that CHC does not exploit favorable scenarios where $\mu$ is highly concentrated around the seed queries of the Voronoi tessellation. It is important to stress that this lower bound is specific to the CHC algorithm and does not extend to all possible algorithms. The proof, presented below, relies on the connection between random polytopes and the so-called floating body of the polytope they approximate, a notion developed in Schütt and Werner (1990).
@@
Lemma E.3 leverages the relationship between the uncovered region $P\setminus P_{n}$ and the wet part of $\nu$, which we now define. For a distribution $\nu$, the wet part of $\nu$ is defined in Bárány et al. (2020) as
@@
To understand this notion intuitively, consider the case where $\nu$ is uniform on the unit ball, and assume that the ball is filled with a volume $t$ of water. The wet part $W^{t}_{\nu}$ represents the “trace” left by the water as the ball rolls on a flat surface, and $w^{\nu}(t)$ is the volume of that wet part. The $\nu$-measure of the wet part is tightly linked to the expected missing volume $\mathbb{E}[1-\nu(P_{n})]$, as shown in e.g. Schütt (1991) and Bárány and Buchta (1993) for the case where $\nu$ is uniform on some convex polytope of $\mathbb{R}^{d}$, and in Bárány et al. (2020) for arbitrary measures in $\mathbb{R}^{d}.$ Such results are sometimes stated in terms of the floating body of $\nu$, which is the relative complement of its wet part in its support, see Schütt and Werner (1990).
@@
By the lower bound of Theorem 1.2 in Bárány et al. (2020):
@@
This lower bound is straightfoward to derive: one can note that $\mathbb{P}(x\notin P_{n})\geq(1-t)^{n}$ for any $x\in W^{t}_{\nu}$, thus $\mathbb{E}[\nu(P\setminus P_{n})]=\int_{0}^{\infty}\mathbb{P}(x\notin P_{n})d\mu(x)\geq(1-t)^{n}w^{\nu}(t)$. Setting $t=1/n$ yields the result (see Section 3.1 in Bárány et al. (2020) for a more detailed proof).
Then, by Corollary 1.2 in Besau et al. (2018), under Assumption 2.3(ii), the asymptotic behavior of $w^{\nu}$ is independent of $\nu$:
@@
    The geometry of each cell $\mathcal{C}_{i}$. This dependency comes from our application of Proposition B.2 (Theorem 2 from Bárány and Buchta (1993)).
@@
Below, we support the claim that it is very difficult to identify the precise dependency in the cell geometry (and in particular in the dimension) induced by Theorem $2$ from Bárány and Buchta (1993), which was not tracked in their original proof. One way to attempt to identify this dependency is to exploit the relationship between random polytopes and the floating body of the polytope they approximate, as already leveraged in the proof of Theorem E.2. For example, Bárány et al. (2020) show in their Theorem 1.2 that if $p_{1},\dots,p_{n}$ are sampled according to some arbitrary $\nu$ on $\mathbb{R}^{d}$ with convex hull $P_{n}$,
@@
Here, the dependency in $(\log{n})/n$ in the wet part measure is necessary because $\nu$ is arbitrary, but for more specific distributions (e.g. log-concave), this dependency may potentially be improved to $1/n$, as discussed in Section 2.4 of Bárány et al. (2020). Since $\mathbb{E}[1-\nu(P_{n})]$ is precisely the measure of the region uncovered by $P_{n}$, we may attempt to upper bound $w^{\nu}(t)$ for fixed $t$ to obtain a fully explicit upper bound on our quantity of interest. This is not too difficult when $\nu$ is uniformly distributed on the hypercube $\mathcal{I}^{d}$, as we show below.
@@
Let $v(x)=\min\{\lambda(\mathcal{I}^{d}\cap H),x\in H\text{ and $H$ is a halfspace}\}$, and $u(x)=\lambda(\mathcal{I}^{d}\setminus(2x-\mathcal{I}^{d}))$. This is the volume of a Macbeath region, see e.g. Bárány (2008), where they recall in their Lemma 4.3 that $u(x)\leq 2v(x)$. Note also that the wet part for the uniform distribution on $\mathcal{I}^{d}$ is $\{x\in K,v(x)\leq t\lambda(\mathcal{I}^{d})\}.$ Then